OpenStudy (zzelinski):

What is point slope form and how can it be used to find the equation of a line?

sam (.sam.):

Point slope \(y-y_1=m(x-x_1)\) Where \((x_1,y_1)\) and slope = m

OpenStudy (zzelinski):

thanks!

OpenStudy (anonymous):

For example, if you want to find an equation of the line passing by the points (1,1) et (3,-1) : First, we can write \[y=mx+b\] where m is the slope which can been found by : \[m=\frac{y_2-y_1}{x_2-x_1}=\frac{2-1}{-1-1}=-1.\] Then, for finding b, we substitute with a point from the line, for example (1,1), so : \[1=-1\times 1+b\] and hence : \[b=-2.\] So the equation of the line is : \[y=-x-2.\]

OpenStudy (zzelinski):

thanks @Noura11 that helped a lot!